OpenStudy (anonymous):

Hello every body, those who are willing to learn Electrical Engineering from basic level, Kinda join me. I will be posting content.

OpenStudy (kenljw):

V =I R P = I V P = I^2 R I = V/R P = V^2/R V volts across a component of delivered by source I amperes, current through a component of delivered by source R resistance in Ohms P watts, power comsumed by compoent or deliverd by source These are very basic equations and units in electrical/electronic courses

rvc (rvc):

basic is the ohm's law :)

OpenStudy (kenljw):

Yes, the first equation I gave, then further developement.

rvc (rvc):

then comes V rms I rms in DC analysis we have mesh,nodal,superposition,thevinins....

OpenStudy (kenljw):

RMS is an abreviation for root mean square in reverse order, it essential represents a signal power corresponding to a DC equivalent, ie effective power. in DC analysis we have mesh,nodal,superposition,thevinins.... These are common means of analysis of an electric/electronic circuits and are used in DC, AC, and signal analysis.